Fermi National Laboratory


Accelerator Update

Wednesday April 28
The day shift began with Operations monitoring the Recycler stash, with the Collider Detector at Fermilab (CDF) and D-zero (D0) experiments in access, with the Antiproton Source (Pbar) stacking, with cryo system technicians working on the Tevatron (TeV) sector A1 cold compressor, and with MiniBooNE and with the Meson MTest experiments taking beam.

The Main Injector (MI) ramp tripped off at 9:30 AM due to an overcurrent indication. Operators could not reset the system and called experts. An expert replaced a bad voltage monitor card, which restored the ramp about thirty minutes later.

At 9:41 AM, operators prepared the TeV for access.

Pbar resumed stacking at 10:03 AM.

Cryo system techs reported finishing the cold compressor work at 10:06 AM.

CDF came out of access at 10:16 AM.

Operations put a stop to Pbar stacking at 11:13 AM due to an emittance problem. Operators resolved the problem and resumed stacking at 11:37 AM.

The TeV access was completed by 11:25 AM.

Operations put the TeV through a dry squeeze at 1:02 PM.

A Linac expert began a Preaccelerator tune-up at 1:14 PM.

Operations put the TeV through a wet squeeze at 2:06 PM and then into shot setup at 2:30 PM. The protons had been loaded by 3:52 PM, but when operators began loading the antiprotons about five bunches of protons got knocked out. Operators put the shot setup on hold while they investigated.

Operations aborted the TeV beam at 4:54 PM.

At 4:55 PM, Operation began running the TeV through several dry squeezes.

At 7:22 PM, TeV experts reported that that they traced the ramp problem to a low level RF module not cogging even though all the readbacks said they were. Experts wanted to repeat the ramp study to see if they could repeat the failure.

Operations put the TeV back into shot setup and, after resolving many system problems, began loading protons at 7:53 PM. By 8:21 PM, operators had finished loading the final antiprotons. At 8:44 PM, Operations established store 3456 with an initial luminosity of 49.86E30. Pbar resumed stacking.

Operations reported at 9:58 PM that both CDF and D0 were off for about thirty minutes due to beam problems, which operators resolved by adjusting the TeV collimators.

At 10 PM, a Booster RF station (BRF18) tripped off. It reset okay.

At 11:08 PM, Operations reported having problems with the Booster beam. Operators disabled beam to MiniBooNE, MTest, and MI. Operators resolved a notcher problem thirty minutes later and resumed sending beam.

Thursday April 29
The midnight shift began with Operations monitoring store 3456 and the stash, with Pbar stacking, with a MI expert conducting studies, and with MiniBooNE and MTest taking beam.

At 12:42 AM, MI began aborting after every DC ramp event. Operators investigated and had beam restored by 1:10 AM.

Operations reported at 2:18 AM that a Booster RF station (BRF13) began regularly tripping off.

At 2:21 AM, a Linac RF station (LRF3) tripped off for the first of several times. Operators had the station back up by 2:49 AM.

Pbar resumed stacking at 2:38 AM.

Two Booster RF stations (BRF13 & 16) began tripping off regularly at 3:03 AM. Operations bypassed the stations and called in a RF tech to make repairs.

At 4:13 AM, Operations turned off Switchyard (SY) beam and put the Recycler into shot setup.

The RF tech completed his repairs on BRF13 & 16 at 4:41 AM.

Operations completed stashing antiprotons and resumed stacking at 5:19 AM.

Operations reestablished beam to MTest at 5:24 AM.

Two Booster RF stations (BRF7 & 8) tripped off at 7:46 AM due to a vacuum burst.

At 7:49 AM, MI aborted again due to an overcurrent indication. An expert was called and came in, but the MI beam resumed ten minutes later without the expert doing anything.

Operations disabled the MTest beam at the experimenter's request at 8:42 AM.

A Linac RF station (LRF3) tripped off at 8:47 AM. Operators and experts investigated and had the station back up by 9:15 AM.

Pbar resumed stacking at 9:45 AM.

Operations turned off Booster, Linac, the Preaccelerator, and MiniBooNE at 10:11 AM for a Booster access.

Recycler experts began a two-hour study period at 10:39 AM.

Linac experts changed out the H- Source's hydrogen bottle at 10:42 AM.

The Booster and Linac accesses ended and Operations had beam to the Accumulator by 2:46 PM.

Operations established beam to MiniBooNE at 3:26 PM.

A TeV RF station (TRF7) tripped off at 3:32 PM due to an upstream water pump failure. Techs will replace this after the present store gets terminated.

A CDF experimenter informed Operations at 3:42 PM that their silicon detector was off due to a cooling problem.

At 6:05 PM, an Accumulator septa tripped off due to an overcurrent. It tripped off again thirty minutes later and then again at 7:28 PM.

A Pbar expert began an hour-long study at 7:19 PM.

Operations terminated store 3456 at 11 PM and then made the D0 collision hall ready for access.

Friday April 30
The midnight shift began with Operations monitoring the stash, with D0 experimenters in access, with TeV RF repairs in progress, with Pbar stacking, and with MiniBooNE taking beam.

D0 came out of access at 12:18 AM.

Operations put the TeV into shot setup at 1:07 AM and at 2:54 AM established store 3458. Its initial luminosity was 52.0E30. Pbar resumed stacking.

At 5:45 AM, operations put the Recycler into shot setup. By 6:21, the Recycler had stashed two transfers of antiprotons. Pbar resumed stacking at 6:41 AM.

Plans for Today and the Weekend
The plans for today and this weekend are to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/30/2004   email Fermilab